OpenStudy (anonymous):

Solve the following equation. One point is given for showing steps and one point given for the correct solution. (5x-1) / 2 + (-3x-1) / 4 = 9/2

OpenStudy (anonymous):

multiply the entire equation by 4

OpenStudy (anonymous):

2 (5x - 1) + - 3x - 1 = 18 10x - 2 - 3x - 1 = 18 7x - 3 = 18 7x = 21 x = 3